OpenStudy (anonymous):

You are given seven cards numbered 1 through 7. You shuffle the cards and deal them out on a table. What is the probability that the cards will be dealt in order: 1, 2, 3, 4, 5, 6, 7?

Parth (parthkohli):

This is actually pretty easy! P(Getting 1) = 1/7 P(Getting 2) = 1/6 (because we have a card less) P(Getting 3) = 1/5 P(4) = 1/4 P(5) = 1/3 P(6) = 1/2 P(7) = 1

Parth (parthkohli):

Now multiply all the P's.